OpenStudy (anonymous):

Write the ratio in lowest terms with whole numbers in the numerator and denominator. 10.15 hr to 8.12 hr (Show steps please & thank you!)

OpenStudy (anonymous):

multiply by 100 and simplify completely

OpenStudy (anonymous):

multiply by 100 both the numerator and the denominator and simplify completely

OpenStudy (anonymous):

How do I get them to lowest terms? Prime factorization?

OpenStudy (vishal_kothari):

http://drrobinson.com/TCC/0002SP07/5-1%20&%205-2%20Ratios%20&%20Rates.pdf exercise no. 1

OpenStudy (anonymous):

rules of divisibility or prime factorization

OpenStudy (anonymous):

helps to know that \[1015=5\times 7\times 29\] and \[812 = 2^2\times 7\times 29\] so that \[\frac{1015}{812}=\frac{5}{7}\]

OpenStudy (anonymous):

oop \[\frac{1015}{812}=\frac{5}{4}\] looks better

OpenStudy (anonymous):

http://www.youtube.com/watch?v=nVbv1mEiHx8

OpenStudy (anonymous):

@satellite wish Lynn could do it ! do you think your method helped her understand ?

OpenStudy (anonymous):

Wait.. why is it 5/4 now? :|

OpenStudy (anonymous):

I got 5/7.. at first.

OpenStudy (anonymous):

...Because 2^2 = 4?

OpenStudy (anonymous):

I think I got it.

OpenStudy (anonymous):

yes that is why you have a choice, either factor both numbers and cancel all common factors, or cancel on piece at a time. but in this case it is unlikely that you would recognize that both numbes are divisible by 29 unless you started out slow. or divisible by 7 for that matter. you would just have to check

OpenStudy (anonymous):

for example 1015 is clearly divisible by 5 so divide by 5 and get \[1015=5\times 203\] but the only way you are going to know that 203 is divisible by 7 is to try it then you will see that \[203=7\times 29\] and since 29 is prime you are done

OpenStudy (anonymous):

Thank you :)